OpenStudy (anonymous):

8 + 3 [3-(-1)^6] help me simplify please.

OpenStudy (anonymous):

- -2.3

OpenStudy (anonymous):

(-1)^6=1 3-1=2 3*2=6 8+6=14 i hope i find out the problem i n correct way

OpenStudy (anonymous):

It was, thanks

OpenStudy (anonymous):

ur welcome